# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ II Problem Formulation
We say that dynamics (3) (or the state $\bm{x}^{\mathsf{N}}$) undergo perpetual fluctuations if there exists some non-singleton set $\bm{\mathcal{Y}}^{\mathsf{N}}\subseteq\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$, the fluctuation set, such that for every state $\bm{y}^{\mathsf{N}}\in\bm{\mathcal{Y}}^{\mathsf{N}}$ and $K>0$, there exists some time $T>K$ such that $\bm{x}^{\mathsf{N}}(T)=\bm{y}^{\mathsf{N}}$. The length of the perpetual fluctuation set $\bm{\mathcal{Y}}^{\mathsf{N}}$ is defined as $\max_{\bm{y}^{\mathsf{N}}_{i},\bm{y}^{\mathsf{N}}_{j}\in\bm{\mathcal{Y}}^{\mathsf{N}}}|\bm{y}^{\mathsf{N}}_{i}-\bm{y}^{\mathsf{N}}_{j}|$.
A minimal positively invariant set is a non-empty subset of the state space such that once the population state enters the set, it remains there afterward under any activation sequence and visits all states contained in the set with a positive probability, resulting in perpetual fluctuations. A positively invariant set is minimal if it does not contain a non-empty positively invariant subset rather than itself.
The dynamics of a finite mixed population consisting of both coordinating and anticoordinating subpopulations may undergo perpetual fluctuations [18], the characterization of which appears to be challenging and remain unsolved. Could the asymptotic behaviour of discrete populations be revealed by investigating the associated mean dynamics, as the population size approaches infinity?
Figure 1: The long-term fluctuations of the population proportion of strategy-$\mathtt{A}$ players for varying population sizes. The circles represent the upper bounds of the invariant sets and the crosses show the lower bounds of the invariant sets. As the population size increases the lower and upper bounds get closer.
@@
Consider a population of $\mathsf{N}$ agents stratified into one anticoordinating subpopulation and five coordinating subpopulations. The thresholds and population proportions of the subpopulations are respectively as follows: $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau^{\prime}_{5}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau^{\prime}_{4}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau^{\prime}_{3}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau^{\prime}_{2}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tau^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.885},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.89},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.604},,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.481},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.444},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.21})}$ $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{\prime}_{5}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{\prime}_{4}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{\prime}_{3}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{\prime}_{2}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}12/30},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3/30},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3/30},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3/30},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1/30},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}8/30}})$. For each value of $\mathsf{N}=30,60,120,\ldots,3840$, the population state was simulated 100 times with different random activation sequences and initial conditions. For each value of $\mathsf{N}$, we considered $30\mathsf{N}$ steps. For each run, the upper and lower bounds of the population proportion of strategy $\mathtt{A}$ players were recorded. The maximum upper and minimum lower bounds for each population size $\mathsf{N}$ is depicted in Figure 1. As population size grows, the lower and upper bounds of the fluctuations in the proportion of strategy-$\mathtt{A}$ players get closer.
@@
Example 2 gives us an intuition that the length of fluctuations in the population proportion of $\mathtt{A}$-players decreases as population size grows. Does this observation extend to every mixed population of coordinators and anticoordinators as the population size approaches infinity i.e., $\mathsf{N}\rightarrow\infty$? (Fig 2). The available results in stochastic approximation theory provide the basis for investigating this question.
@@
Figure 2: How does the asymptotic behaviour of the finite population of interacting agents (3) evolve as the population size approaches infinity? An exclusive finite population of coordinators and an exclusive finite population of anticoordinators each equilibrates in the long-run, [14, 15]. In the latter case, for some specific configurations, the population may admit a cycle of length 2. The proportion of $\mathtt{A}$-players in a mixed finite population may undergo perpetual fluctuations [18].
@@ III Background
We present the following definitions laying the basis for introducing Theorem 1 which connects the asymptotic behaviour of the finite populations and their associated mean dynamics. More definitions are provided in the appendix.
@@
How to approximate the realizations of a collection of Markov chains $\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}$ for a sequence of $\epsilon>0$ with solutions of a differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$? The following definition establishes the basis for doing so.
@@
[37] Consider the good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ over the convex and compact state space $\bm{\mathcal{X}}_{0}$. For a sequence of positive values of scalar $\epsilon$ approaching 00, let $\textbf{\em U}^{\epsilon}=\langle\textbf{\em U}^{\epsilon}_{k}\rangle_{k}$ be a sequence of $\mathbb{R}^{n}$-valued random variables and $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ be a family of set-valued maps on $\mathbb{R}^{n}$. We say that $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}\rangle_{\epsilon>0}$ is a family of generalized stochastic approximation processes (or a GSAP) for the differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ if the following conditions are satisfied:
@@
    ${\bm{\mathcal{V}}^{\epsilon}(\bm{x})\subset\{\bm{z}\in\mathbb{R}^{n}\!\mid\!\exists\bm{y}:\!|\bm{x}-\bm{y}|<\delta,\!\inf_{\bm{v}\!\in\bm{\mathcal{V}}(\bm{y})}|\bm{z}-\bm{v}|<\delta\},}$
@@
The GSAPs can be regarded as perturbed solutions of the differential inclusions. The first condition ensures that the perturbed solutions are within the same set as the unperturbed solutions are. The second condition defines $\langle\mathbf{X}^{\epsilon}_{k}\rangle$ as a sequence of elements which are recursively updated. At each index $k$, the value of $\mathbf{X}^{\epsilon}_{k+1}$ is characterized based on those of $\mathbf{X}^{\epsilon}_{k}$, $\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$, and the random variable $\textbf{\em U}^{\epsilon}_{k+1}$, which perturbs the value of $\mathbf{X}^{\epsilon}_{k+1}$. The third condition ensures that the collection $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ approaches to the vicinity of $\bm{\mathcal{V}}$ as $\epsilon$ approaches zero. More specifically, the third condition is indeed a limit condition, which requires that for an arbitrarily small $\delta>0$ there exists $\epsilon_{0}>0$ such that for each $\epsilon<\epsilon_{0}$ in the sequence $\langle\epsilon\rangle$, the values belonging to the set valued map $\bm{\mathcal{V}}^{\epsilon}(\bm{x})$ are within a $\delta$-neighborhood of the image of the $\delta$-neighborhood of $\bm{x}$. Well-behaved random variables such as sub-Gaussian random variables and some classes of random variables with finite higher moments satisfy the last condition.
In [37], Roth and Sandholm studied the limiting stationary distributions of these GSAPs and showed that if for each $\epsilon>0$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}$ is a Markov chain, their stationary measures as $\epsilon$ approaches zero will be concentrated on $\mathrm{BC}(\bm{\Phi})$. The following theorem is a result of 1 and was implicitly used in the proof of [37, Theorem 4.11].
@@
For a vanishing sequence $\langle\epsilon_{k}\rangle=\epsilon_{0},\epsilon_{1},\ldots,$ let $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}\rangle_{\epsilon}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon_{n}$, $n\in\mathbb{Z}_{\geq 0}$, $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle_{k}$ is a Markov chain and $\bm{\mu}^{\epsilon_{n}}$ is an invariant probability measure of $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle_{k}$. We then have $\lim_{n\to\infty}\bm{\mu}^{\epsilon_{n}}(\mathcal{O})=1$, where $\mathcal{O}$ is any open set containing the Birkhoff center of the dynamical system $\bm{\Phi}$ induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
IV The link to the semicontinuous dynamics
Our aim is to determine the asymptotic behaviour of the population dynamics (3) as the population size $\mathsf{N}$ approaches infinity. Theorem 1 connects the asymptotic behaviour of the GSAPs to the steady-state behaviour of their associated differential inclusion. Hence, if one shows that firstly, dynamics (3) define a Markov chain and secondly, the collection of the Markov chains indexed by the population size is a GSAP for a good upper semicontinuous differential inclusion, then the results of Theorem 1 can be applied, meaning that the asymptotic behaviour of the population dynamics can be revealed by analyzing the differential inclusion.
In view of (3), the next state of the population dynamics only depends on the current population state and the adopted strategy of and the subpopulation associated with the current active agent. As a result, the sequence of $\langle\bm{x}^{\mathsf{N}}(k)\rangle_{k}$ is a realization of a Markov chain.
@@ IV The link to the semicontinuous dynamics
The sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle_{k}$ is a realization of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0)$ and transition probabilities
$\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}=\begin{cases}({\rho}_{p}-x^{\mathsf{N}}_{p})(2-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{*}}(p,\bm{x}^{\mathsf{N}}))&\text{if }\bm{y}^{\mathsf{N}}=\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ x^{\mathsf{N}}_{p}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{*}}(p,\bm{x}^{\mathsf{N}})-1)&\text{if }\bm{y}^{\mathsf{N}}=-\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\big{(}}x^{\mathsf{N}}_{p}(2-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{*}}(p,\bm{x}^{\mathsf{N}}))&\text{if }\bm{y}^{\mathsf{N}}=\bm{x}^{\mathsf{N}},\\ \hskip 33.0pt+({\rho}_{p}-x^{\mathsf{N}}_{p})({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{*}}(p,\bm{x}^{\mathsf{N}})&-1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\big{)}}\\ 0&\text{otherwise},\end{cases}$
where $\mathbf{e}_{p}$ is the $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional vector with all elements equal to zero except for its $p$th element, which is equal to 1.
@@
The population dynamics Markov chain is the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0)$ and transition probabilities (4).
Remark 2.
We increase the population size $\mathsf{N}$ in a way that the population structure, i.e., the population proportions of the subpopulations $\bm{\rho}$ remains unchanged. Hence, the sequence $\langle\mathsf{N}\rangle$ according to which the population size approaches infinity should satisfy $\mathsf{N}\bm{\rho}\in\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$. We denote the set of such valid population sizes by $\mathcal{N}$.
Now, to apply Theorem 1, we need to show that the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for a good upper semicontinuous differential inclusion and hence the support of the limit point of its invariant probability measures is determined by the steady-state behavior of that differential inclusion. We claim that $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the following differential inclusion:
@@
The semicontinuous population dynamics is defined by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x}),$ where $\bm{\mathcal{V}}:\bm{\mathcal{X}}_{ss}\rightarrow 2^{\bm{\mathcal{X}}_{ss}}$, and for $p\leq\mathsf{p}$

### v3

@@ II Problem Formulation
A fluctuation set for the dynamics (3) is a non-singleton set $\bm{\mathcal{Y}}^{\mathsf{N}}\subseteq\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$ such that if $\bm{x}^{\mathsf{N}}(k)\in\bm{\mathcal{Y}}^{\mathsf{N}}$, for some $k$, then (i) $\bm{x}^{\mathsf{N}}(k^{\prime})\in\bm{\mathcal{Y}}^{\mathsf{N}}$ for all $k^{\prime}>k$ and (ii) for every state $\bm{y}^{\mathsf{N}}\in\bm{\mathcal{Y}}^{\mathsf{N}}$ and every $K>0$ and under every persistent realizations of $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$ , there exists some $T>K$ such that $\bm{x}^{\mathsf{N}}(k+T)=\bm{y}^{\mathsf{N}}$. The amplitude of the perpetual fluctuation set $\bm{\mathcal{Y}}^{\mathsf{N}}$ is defined as $\max_{\bm{y}^{\mathsf{N}},\bm{z}^{\mathsf{N}}\in\bm{\mathcal{Y}}^{\mathsf{N}}}|\bm{1}^{\top}\bm{y}^{\mathsf{N}}-\bm{1}^{\top}\bm{z}^{\mathsf{N}}|.$ Starting from a given initial condition and under a specified realization of $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$, we say that the dynamics (3) undergo perpetual fluctuations if $\bm{x}^{\mathsf{N}}(k)\in\bm{\mathcal{Y}}^{\mathsf{N}}$ for some fluctuation set $\bm{\mathcal{Y}}^{\mathsf{N}}$ and finite $k\geq 0$.
The dynamics of a finite mixed population consisting of both coordinating and anticoordinating subpopulations may undergo perpetual fluctuations [19], the characterization of which appears to be challenging and remain unsolved. Could the asymptotic behavior of discrete populations be revealed by investigating the associated mean dynamics, as the population size approaches infinity?
Figure 1: The long-term fluctuations of the population proportion of strategy-$\mathtt{A}$ players for varying population sizes. The circles and crosses represent the maximum and minimum of the population proportion of $\mathtt{A}$-players, respectively. As the population size increases, these values get closer.
@@
Consider a population of $\mathsf{N}$ agents stratified into one anticoordinating and five coordinating subpopulations. The thresholds and population proportions of the subpopulations are respectively as follows: $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau_{1}},{\tau^{\prime}_{5}},{\tau^{\prime}_{4}},{\tau^{\prime}_{3}},{\tau^{\prime}_{2}},{\tau^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.885},{0.89},{0.604},,{0.481},{0.444},{0.21})}$ $\smash{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}},{\rho^{\prime}_{5}},{\rho^{\prime}_{4}},{\rho^{\prime}_{3}},{\rho^{\prime}_{2}},{\rho^{\prime}_{1}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}12/30},{3/30},{3/30},{3/30},{1/30},{8/30}})$. For each value of $\mathsf{N}=30,60,120,\ldots,3840$, the population state was simulated 100 times with different random activation sequences and initial conditions. For each value of $\mathsf{N}$, we considered $30\mathsf{N}$ steps. For each run, the maximum and minimum of the population proportion of strategy $\mathtt{A}$ players were recorded. These values for each population size $\mathsf{N}$ are depicted in Figure 1. As population size grows, the maximum and minimum of the proportion of strategy-$\mathtt{A}$ players get closer and eventually merge.
@@
Example 2 gives us an intuition that the amplitude of fluctuations in the population proportion of $\mathtt{A}$-players decreases as population size grows. Does this observation extend to every mixed population of coordinators and anticoordinators as the population size approaches infinity i.e., $\mathsf{N}\rightarrow\infty$? (Fig 2). The available results in stochastic approximation theory provide the basis for investigating this question.
@@
Figure 2: How does the asymptotic behavior of the finite population of interacting agents (3) evolve as the population size approaches infinity? An exclusive finite population of coordinators and an exclusive finite population of anticoordinators each equilibrates in the long-run, [15, 16]. In the latter case, for some specific configurations, the population may admit a two-state cycle. The proportion of $\mathtt{A}$-players in a mixed finite population may undergo perpetual fluctuations [19].
@@ III Background
We present the following definitions laying the basis for introducing Theorem 1, which connects the asymptotic behavior of the finite populations and their associated mean dynamics. Please see the appendix for more details.
A differential inclusion is defined by
$\dot{\bm{x}}(t)\in\bm{\mathcal{V}}\big{(}\bm{x}(t)\big{)},$ (4)
where $\bm{\mathcal{V}}\subset\mathbb{R}^{n}$ is a set-valued map specifying a set of evolution at each point $\bm{x}\in\mathbb{R}^{n}$.
@@
How to approximate the realizations of a collection of Markov chains $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ for a sequence of $\epsilon>0$ with solutions of a differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$? The following definition establishes the basis for doing so.
@@
[38] Consider the good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ over the convex and compact state space $\bm{\mathcal{X}}_{0}$. For a sequence of positive values of scalar $\epsilon$ approaching 00, let $\textbf{\em U}^{\epsilon}=\langle\textbf{\em U}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ be a sequence of $\mathbb{R}^{n}$-valued random variables and $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ be a family of set-valued maps on $\mathbb{R}^{n}$. We say that $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\epsilon>0}$ is a family of generalized stochastic approximation processes (or a GSAP) for the differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ if the following conditions are satisfied:
@@
    $\bm{\mathcal{V}}^{\epsilon}(\bm{x})\subset\{\bm{z}$ $\in\mathbb{R}^{n}\mid\exists\bm{y}:|\bm{x}-\bm{y}|<\delta,$ (5)
    $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta\},$
@@
The GSAPs can be regarded as perturbed solutions of the differential inclusions. The first condition ensures that the perturbed solutions are within the same set as the unperturbed solutions are. The second condition defines $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$ as a sequence of elements which are recursively updated. At each index $k$, the value of $\mathbf{X}^{\epsilon}_{k+1}$ is characterized based on those of $\mathbf{X}^{\epsilon}_{k}$, $\bm{\mathcal{V}}^{\epsilon}(\mathbf{X}^{\epsilon}_{k})$, and the random variable $\textbf{\em U}^{\epsilon}_{k+1}$. The third condition ensures that the collection $\langle\bm{\mathcal{V}}^{\epsilon}\rangle$ approaches to the vicinity of $\bm{\mathcal{V}}$ as $\epsilon$ approaches zero. More specifically, the third condition is indeed a limit condition, which requires that for an arbitrarily small $\delta>0$ there exists $\epsilon_{0}>0$ such that for each $\epsilon<\epsilon_{0}$ in the sequence $\langle\epsilon\rangle$, the values belonging to the set valued map $\bm{\mathcal{V}}^{\epsilon}(\bm{x})$ are within a $\delta$-neighborhood of the image of the $\delta$-neighborhood of $\bm{x}$. Well-behaved random variables, such as sub-Gaussian random variables and some classes of random variables with finite higher moments satisfy the last condition.
In [38], Roth and Sandholm studied the limiting stationary distributions of these GSAPs and showed that if for each $\epsilon>0$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle$ is a Markov chain, their stationary measures as $\epsilon$ approaches zero will be concentrated on $\mathrm{BC}(\bm{\Phi})$. The following theorem is a result of A1 and was implicitly used in the proof of [38, Theorem 4.11].
@@
For a vanishing sequence $\langle\epsilon_{n}\rangle_{n=0}^{\infty}$ let $\langle\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{n}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon_{n}$, $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ is a Markov chain and ${\mu}^{\epsilon_{n}}$ is an invariant probability measure of $\langle\mathbf{X}^{\epsilon_{n}}_{k}\rangle{{{}_{k=0}^{\infty}}}$. We then have $\lim_{n\to\infty}{\mu}^{\epsilon_{n}}(\bm{\mathcal{O}})=1$, where $\bm{\mathcal{O}}$ is any open set containing the Birkhoff center of the dynamical system $\bm{\Phi}$ induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
IV The link to the continuous-time dynamics
Our aim is to determine the asymptotic behavior of the population dynamics (3) as the population size $\mathsf{N}$ approaches infinity. Theorem 1 connects the asymptotic behavior of the GSAPs to the steady-state behavior of their associated differential inclusion. Hence, if one shows that, firstly, dynamics (3) define a Markov chain, and secondly, the collection of the Markov chains indexed by the population size is a GSAP for a good upper semicontinuous differential inclusion, then the results of Theorem 1 can be applied, meaning that the asymptotic behavior of the population dynamics can be revealed by analyzing the differential inclusion.
In view of (3), the next state of the population dynamics only depends on the current population state and the adopted strategy of and the subpopulation associated with the current active agent. As a result, the sequence of $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{{}_{k=0}^{\infty}}}$ is a realization of a Markov chain.
@@ IV The link to the continuous-time dynamics
The sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle{{{}_{k=0}^{\infty}}}$ is a realization of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0)$ and transition probabilities
$\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}=$
$\begin{cases}({\rho}_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))&\text{if }\bm{y}^{\mathsf{N}}=\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=-\frac{1}{\mathsf{N}}\mathbf{e}_{p}+\bm{x}^{\mathsf{N}},\\ {1-}\displaystyle\sum_{p=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\big{(}x^{\mathsf{N}}_{p}({s^{*}}(p,\bm{x}^{\mathsf{N}},{1})-1)&\text{if }\bm{y}^{\mathsf{N}}=\bm{x}^{\mathsf{N}},\\ \hskip 27.0pt+(\rho_{p}-x^{\mathsf{N}}_{p})(2-{s^{*}}(p,\bm{x}^{\mathsf{N}},{2}))\big{)}\\ 0&\text{otherwise},\end{cases}$ (6)
where $\mathbf{e}_{p}$ is the $p^{\text{th}}$ standard basis vector in $\mathbb{R}^{(\mathsf{p}+\mathsf{p}^{\prime})}$.
@@
The population dynamics Markov chain is the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}$ with the state space $\bm{\mathcal{X}}_{ss}\cap\frac{1}{\mathsf{N}}\mathbb{Z}^{{\mathsf{p}+\mathsf{p}^{\prime}}}$, initial state $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{0}=\bm{x}^{\mathsf{N}}(0),$ and transition probabilities (6).
We increase the population size $\mathsf{N}$ in a way that the population structure, i.e., the population proportions of the subpopulations $\bm{\rho}$ remains unchanged. Hence, the sequence according to which the population size approaches infinity should satisfy $\mathsf{N}\bm{\rho}\in\mathbb{Z}^{\mathsf{p}+\mathsf{p}^{\prime}}$. We denote the set of such valid population sizes by $\mathcal{N}$.
Now, to apply Theorem 1, we need to show that the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for a good upper semicontinuous differential inclusion and hence the support of the limit point of its invariant probability measures is determined by the steady-state behavior of that differential inclusion. We claim that $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the following differential inclusion:
@@
The continuous-time population dynamics are defined by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$ where $\bm{\mathcal{V}}:\bm{\mathcal{X}}_{ss}\rightrightarrows\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$, for $p\leq\mathsf{p}$